the sum of two consecutive oddintegers is 244. what is the smaller integer?
Answer by checkley71(8403)

X+(X+2)=244 OR 2X+2=244 OR 2X=244-2 OR 2X=242 OR X=242/2 OR X=121 IS THE SMALLER INTERGER & X+2=123
121+123=244